# ReZero: Boosting MCTS-based Algorithms by Backward-view and Entire-buffer Reanalyze

### v4

@@ Appendix A Proof Materials
Lemma A.1.
[…]
$\kappa=\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{t}+\sqrt{\frac{a}{t^{2}}}\geq\varepsilon\}$ (9)
, then it holds that $E[\kappa]\leq 1+\frac{2\sqrt{a}}{\varepsilon}+\frac{C^{2}}{\varepsilon^{2}}$.
@@
$E[\kappa]$ $\leq u+\sum_{t=\lceil u\rceil}^{n}\mathbb{P}(\hat{\mu}_{t}+\sqrt{\frac{a}{t^{2}}}\geq\varepsilon)$ (10)
@@
We first present the upper bound of $\mathbb{E}[T_{i}(n)]$ for using the Equation 3 in AlphaZero. A slight adjustment to this proof yields the conclusion of Theorem 1. Denote $T_{i}(k)$ as the times that arm $i$ has been chosen until time $k$, $\hat{\mu}_{is}$ as the average of the first $s$ samples of arm $i$, which satisfies the concentration assumption in Equation 2, and $\hat{\mu}_{i}(k)=\hat{\mu}_{iT_{i}(k)}$. Then we have:
@@
for 16, we have
$E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (18)
@@
$\leq 1+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}$ (24)
Notes: In 19, since we assume $P_{1}\geq\mu_{1}$, the probability of the term $s=0$ would be 00. Thus, we can discard it. If this assumption doesn’t hold, we can choose to accumulate $t$ starting from a larger $t_{0}$(which satisfies $P_{1}\sqrt{t_{0}-1}>\mu_{1}$ as mentioned in the article). Starting the summation from such a $t_{0}$ ensures all terms of $s=0$ can still be discarded, and all add terms that $t\leq t_{0}$ can be bounded to $1$. This only changes the constant term and won’t affect the growth rate of regret. From 21 to 22, we just need to expand the quadratic term and do some simple inequality scaling. From 22 to 23, we need to notice that $\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{C^{2}})$ is a geometric sequence and scale it to $\frac{C^{2}}{\varepsilon^{2}}$. From 23 to 24, we use the inequality $\sum_{t=2}^{n}\frac{1}{e^{a\sqrt{t-1}}}\leq\int_{1}^{\infty}\frac{1}{e^{a\sqrt{t-1}}}$.
for 17, we have
$E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (25)
$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\sqrt{\frac{n-1}{(1+T_{i}(t-1))^{2}}}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (26)
$\leq 1+E[\sum_{s=1}^{n-1}\mathbb{I}\{\hat{\mu}_{is}+P_{i}\sqrt{\frac{n-1}{(1+s)^{2}}}\geq\mu_{1}-\varepsilon\}]$ (27)
$\leq 1+E[\sum_{s=1}^{n}\mathbb{I}\{\hat{\mu}_{is}-\mu_{i}+P_{i}\sqrt{\frac{n-1}{s^{2}}}\geq\Delta_{i}-\varepsilon\}]$ (28)
with Lemma A.1, we can have
@@
so
$E[T_{i}(n)]\leq 3+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}$ (30)
The proof of theorem 1 can be obtained by making slight modifications. In case we have $\hat{\mu}^{*}_{n}=\frac{1}{n}\sum_{t=1}^{n}\hat{\mu}_{t}$ where $\hat{\mu}_{t}$ is sampled from the optimal arm in advance,
$E[T_{i}(n)]=E[\sum_{t=1}^{n}\mathbb{I}\{A_{t}=i\}]$ (31)
$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1n}\leq\mu_{1}-\varepsilon\}]$ (32)
$+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (33)
$\leq n\exp{(-\frac{n\varepsilon^{2}}{C^{2}})}+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (34)
@@
In case we have $\hat{\mu}_{ln}=\frac{1}{n}\sum_{t=1}^{n}\hat{\mu}_{t}$ where $\hat{\mu}_{t}$ is sampled from the same reward distribution of arm $l$,
$E[T_{l}(n)]=E[\sum_{t=1}^{n}\mathbb{I}\{A_{t}=l\}]$ (36)
$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (37)
$+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{ln}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (38)
$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (39)
$\leq 1+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (40)
@@
Appendix B Implementation Details
B.1 Environments
@@
B.2 Algorithm Implementation Details
Our algorithm’s implementation is based on the open-source code of LightZero [30]. Given that our proposed theoretical improvements are applicable to any MCTS-based RL method, we have chosen MuZero and EfficientZero as case studies to investigate the practical improvements in time efficiency achieved by integrating the ReZero boosting techniques: just-in-time reanalyze and speedy reanalyze (temporal information reuse).
To ensure an equitable comparison of wall-clock time, all experimental trials were executed on a fixed single worker hardware setting consisting of a single NVIDIA A100 GPU with 30 CPU cores and 120 GiB memory. Besides, we emphasize that to ensure a fair comparison of time efficiency and sample efficiency, the model architecture and hyper-parameters used in the experiments of Section 5 are essentially consistent with the settings in LightZero [30]. For specific hyper-parameters of ReZero-M and MuZero on Atari, please refer to the Table 3. The main different hyper-parameters for the ReZero-M algorithm in the Connect4 and Gomoku environment are set out in Tables 4. In addition to employing an LSTM network with a hidden state dimension of 512 to predict the value prefix [21], all hyperparameters of ReZero-E are essentially identical to those of ReZero-M in Table 3.
@@ B.2 Algorithm Implementation Details
  • collect time: The total time spent by an agent interacting with the environment to gather experience data.[15] can be integrated to speed up. Besides, this design also makes MCTS-based algorithms compatible to existing RL exploration methods like [34].
  • reanalyze time: The time used to reanalyze collected data with the current policy or value function for more accurate learning targets [13].
@@
Currently, we have set collect_max_episode_steps to 10,000 and eval_max_episode_steps to 20,000 to mitigate the impact of anomalously long evaluation episodes on time. In the future, we will consider conducting offline evaluations to avoid the influence of evaluation time on our measurement of time efficiency. Furthermore, the ReZero methodology represents a pure algorithmic enhancement, eliminating the need for supplementary computational resources or additional overhead. This approach is versatile, enabling seamless integration with single-worker serial execution environments as well as multi-worker asynchronous frameworks. The exploration of ReZero’s extensions and its evaluations in a multi-worker [16] paradigm are earmarked for future investigation.
@@
Replay Ratio [32] 0.25
Reanalyze ratio [13] 1
@@
Reward clipping [28] True
Optimizer type Adam [35]
@@
Weight of SSL (self-supervised learning) loss [21] 2
@@
Normalization type Batch Normalization [36]
Priority exponent coefficient [37] 0.6
@@
The scale of supports used in categorical distribution [38] 300
@@
Replay Ratio [32] 0.25
Reanalyze ratio [13] 1
@@
Appendix C Additional Experiments
C.1 Sample-efficiency of ReZero-M
Except for the results reported in main paper for time efficiency, we also shows the detailed sample efficiency results in this section. Figure 6 displays the performance over environment interaction steps of the ReZero-M algorithm compared with the original MuZero algorithm across six representative Atari environments and two board games. We can find that ReZero-M obtained similar sample efficiency than MuZero on the most tasks.

### v5

@@ Appendix A Proof Materials
Lemma A: Let $w_{t}$ be a random variable that satisfies the concentration condition in Equation 2 with zero expectation, $\varepsilon>0$, $a>0$ and
$\kappa=\sum_{t=1}^{n}\mathbb{I}\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}w_{t}}+\sqrt{\frac{a}{t^{2}}}\geq\varepsilon\}$ (9)
then it holds that $E[\kappa]\leq 1+\frac{2\sqrt{a}}{\varepsilon}+\frac{C^{2}}{\varepsilon^{2}}$.
@@
$E[\kappa]$ $\leq u+\sum_{t=\lceil u\rceil}^{n}\mathbb{P}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}w_{t}}+\sqrt{\frac{a}{t^{2}}}\geq\varepsilon)$ (10)
@@
We first present the upper bound of $\mathbb{E}[T_{i}(n)]$ for using the Equation 3 in AlphaZero. A slight adjustment to this proof yields the conclusion of Theorem 1. Denote $T_{i}(k)$ as the number of times that arm $i$ has been chosen until time $k$, $A_{t}$ as the arm selected at time $t$, $\hat{\mu}_{is}$ as the average of the first $s$ samples of arm $i$, $\mu_{i}$ as the limit of $\mathbb{E}[\hat{\mu}_{is}]$, which satisfies the concentration assumption in Equation 2, and $\hat{\mu}_{i}(k)=\hat{\mu}_{iT_{i}(k)}$. Without loss of generality, we assume that arm $1$ is the optimal arm. Then we have:
@@
for Equation 16, we have
$\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (18)
@@
$\leq 1+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{2C^{6}}{\varepsilon^{4}P_{1}^{2}}}$ (24)
Notes: In Equation 19, since we assume $P_{1}\geq\mu_{1}$, the probability of the term $s=0$ would be 00. Thus, we can discard it. If this assumption doesn’t hold, we can choose to accumulate $t$ starting from a larger $t_{0}$(which satisfies $P_{1}\sqrt{t_{0}-1}>\mu_{1}$ as mentioned in the article). Starting the summation from such a $t_{0}$ ensures all terms of $s=0$ can still be discarded, and all add terms that $t\leq t_{0}$ can be bounded to $1$. This only changes the constant term and won’t affect the growth rate of regret. From Equation 21 to Equation 22, we just need to expand the quadratic term and do some simple inequality scaling. From Equation 22 to Equation 23, we need to notice that $\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{C^{2}})$ is a geometric sequence and scale it to $\frac{C^{2}}{\varepsilon^{2}}$. From Equation 23 to Equation 24, we use the inequality $\sum_{t=2}^{n}\frac{1}{e^{a\sqrt{t-1}}}\leq\int_{1}^{\infty}\frac{1}{e^{a\sqrt{t-1}}}$.
And for Equation 17, we have
$\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (25)
$\leq\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\sqrt{\frac{n-1}{(1+T_{i}(t-1))^{2}}}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (26)
$\leq 1+\mathbb{E}[\sum_{s=1}^{n-1}\mathbb{I}\{\hat{\mu}_{is}+P_{i}\sqrt{\frac{n-1}{(1+s)^{2}}}\geq\mu_{1}-\varepsilon\}]$ (27)
$\leq 1+\mathbb{E}[\sum_{s=1}^{n}\mathbb{I}\{\hat{\mu}_{is}-\mu_{i}+P_{i}\sqrt{\frac{n-1}{s^{2}}}\geq\Delta_{i}-\varepsilon\}]$ (28)
with Lemma A, we can have
@@
so we have
$\mathbb{E}[T_{i}(n)]\leq 3+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{2C^{6}}{\varepsilon^{4}P_{1}^{2}}}$ (30)
The proof of theorem 1 can be obtained by making slight modifications. In case we have drawn $n$ samples from the same non-stationary distribution as arm $1$, and the average of these first $n$ samples is $\hat{\mu}_{1}$,
$\mathbb{E}[T_{i}(n)]=\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{A_{t}=i\}]$ (31)
$\leq\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\mu}_{1}}\leq\mu_{1}-\varepsilon\}]$ (32)
$+\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (33)
$\leq n\exp{(-\frac{n\varepsilon^{2}}{C^{2}})}+\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (34)
@@
In case we have drawn $n$ samples from the same non-stationary distribution as arm $l$, and the average of these first $n$ samples is $\hat{\mu}_{l}$,
$\mathbb{E}[T_{l}(n)]=\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{A_{t}=l\}]$ (36)
$\leq\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (37)
$+\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{\hat{\mu}_{l}}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}l}\}]$ (38)
$\leq\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (39)
$\leq 1+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{2C^{6}}{\varepsilon^{4}P_{1}^{2}}}+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (40)
@@
Appendix B MuZero
During the inference phase, the representation model transforms a sequence of the last $l$ observations $o_{t-l:t}$ into a corresponding latent state representation $s_{t}$. The dynamics model processes this latent state alongside an action $a_{t}$, yielding the subsequent latent state $s_{t+1}$ and an estimated reward $r_{t}$. Finally, the prediction model accepts a latent state and produces both the predicted policy $p_{t}$ and the state’s value estimate $v_{t}$. These outputs are instrumental in guiding the agent’s action selection process throughout its MCTS. Lastly the agent selects or samples the best action $a_{t}$ following the searched visit count distribution. During the training phase, given a training sequence $\{o_{t-l:t+K},a_{t+1:t+K},u_{t+1:t+K},\pi_{t+1:t+K},z_{t+1:t+K}\}$ at time $t$ sampled from the replay buffer, where $u_{t+k}$ denotes the actual reward obtained from the environment, $\pi_{t+k}$ represents the target policy obtained through MCTS during the agent-environment interaction, and $z_{t+k}$ is the value target computed using n-step bootstrapping Hessel et al. (2018). The representation model initially converts the sequence of observations $o_{t-l:t}$ into the latent state $s_{t}^{0}$. Subsequently, the dynamic model executes $K$ latent space rollouts based on the sequence of actions $a_{t+1:t+K}$. The latent state derived after the $k$-th rollout is denoted as $s_{t}^{k}$, with the corresponding predicted reward indicated as $r_{t}^{k}$. Upon receiving $s_{t}^{k}$, the prediction model generates a predicted policy $p_{t}^{k}$ and a estimated value $v_{t}^{k}$. The final training loss encompasses three components: the policy loss ($l_{p}$), the value loss ($l_{v}$), and the reward loss ($l_{r}$):
$L_{\text{MuZero}}=$ $\sum_{k=0}^{K}l_{p}(\pi_{t+k},p_{t}^{k})+\sum_{k=0}^{K}l_{v}(z_{t+k},v_{t}^{k})+\sum_{k=1}^{K}l_{r}(u_{t+k},r_{t}^{k})$ (41)
MuZero Reanalyze, as introduced in Schrittwieser et al. (2021), is an advanced iteration of the original MuZero algorithm. This variant enhances the model’s accuracy by conducting a fresh Monte Carlo Tree Search on sampled states with the most recent version of the model, subsequently utilizing the refined policy from this search to update the policy targets. Such reanalysis yields targets of superior quality compared to those obtained during the initial data collection phase. The Schrittwieser et al. (2021) expands upon this approach, formalizing it as a standalone method for policy refinement. This innovation opens avenues for its application in offline settings, where interactions with the environment are not possible.
Appendix C Implementation Details
C.1 Environments
@@
DMControl: This continuous control suite comprises 39 continuous control tasks. Our focus here is to validate the effectiveness of ReZero in the continuous action space. Consequently, we have utilized two representative tasks (ball_in_cup-catch and walker-stand) for illustrative purposes. A comprehensive benchmark for this domain will be included in future versions.
@@
C.2 Algorithm Implementation Details
Our algorithm’s implementation is based on the open-source code of LightZero (Niu et al., 2023). Given that our proposed theoretical improvements are applicable to any MCTS-based RL method, we have chosen MuZero and EfficientZero as case studies to investigate the practical improvements in time efficiency achieved by integrating the ReZero boosting techniques: just-in-time reanalyze and speedy reanalyze (temporal information reuse).
To ensure an equitable comparison of wall-clock time, all experimental trials were executed on a fixed single worker hardware setting consisting of a single NVIDIA A100 GPU with 30 CPU cores and 120 GiB memory. Besides, we emphasize that to ensure a fair comparison of time efficiency and sample efficiency, the model architecture and hyper-parameters used in the experiments of Section 5 are essentially consistent with the settings in LightZero. For specific hyper-parameters of ReZero-M and MuZero on Atari, please refer to the Table 3. The main different hyper-parameters in the DMControl task are set out in Tables 4. The main different hyper-parameters for the ReZero-M algorithm in the Connect4 and Gomoku environment are set out in Tables 5. In addition to employing an LSTM network with a hidden state dimension of 512 to predict the value prefix (Ye et al., 2021), all hyperparameters of ReZero-E are essentially identical to those of ReZero-M in Table 3.
@@ C.2 Algorithm Implementation Details
  • collect time: The total time spent by an agent interacting with the environment to gather experience data.Weng et al. (2022) can be integrated to speed up. Besides, this design also makes MCTS-based algorithms compatible to existing RL exploration methods like Burda et al. (2018).
  • reanalyze time: The time used to reanalyze collected data with the current policy or value function for more accurate learning targets (Schrittwieser et al., 2021).
@@
Currently, we have set collect_max_episode_steps to 10,000 and eval_max_episode_steps to 20,000 to mitigate the impact of anomalously long evaluation episodes on time. In the future, we will consider conducting offline evaluations to avoid the influence of evaluation time on our measurement of time efficiency. Furthermore, the ReZero methodology represents a pure algorithmic enhancement, eliminating the need for supplementary computational resources or additional overhead. This approach is versatile, enabling seamless integration with single-worker serial execution environments as well as multi-worker asynchronous frameworks. The exploration of ReZero’s extensions and its evaluations in a multi-worker (Mei et al., 2023) paradigm are earmarked for future investigation.
@@
Replay Ratio (Schwarzer et al., 2023) 0.25
Reanalyze frequency 1
@@
Reward clipping (Mnih et al., 2013) True
Optimizer type Adam
@@
Weight of SSL (self-supervised learning) loss (Ye et al., 2021) 2
@@
Normalization type Layer Normalization
Priority exponent coefficient (Schaul et al., 2015) 0.6
@@
The scale of supports used in categorical distribution (Pohlen et al., 2018) 300
@@
Replay ratio (Schwarzer et al., 2023) 0.25
Reanalyze frequency 1
Batch size 64
Num of frames stacked 1
Num of frames skip 2
Discount factor 0.997
Length of game segment 8
Use augmentation False
Number of simulations in MCTS (sim) 50
Number of sampled actions (Hubert et al., 2021) 20
Table 4: Key hyperparameters of ReZero-M on two DMControl tasks (ball_in_cup-catch and walker-stand). More experiments about this hyper-parameter will be explored in the future version. Other unmentioned parameters are the same as that in Atari settings.
Hyperparameter Value
Replay ratio (Schwarzer et al., 2023) 0.25
Reanalyze frequency 1
@@
Appendix D Additional Experiments
D.1 ReZero-M
In this section, we provide additional experimental results for ReZero-M. As a supplement to Table 1, Table 6 presents the complete experimental results on the 26 Atari environments.
Env. Name ReZero-M MuZero
Alien $\mathbf{1.6}\scriptstyle{\pm 0.2}$ $8.6\scriptstyle{\pm 0.4}$
Amidar $\mathbf{1.5}\scriptstyle{\pm 0.2}$ $8.1\scriptstyle{\pm 0.3}$
Assault $\mathbf{1.5}\scriptstyle{\pm 0.1}$ $7.5\scriptstyle{\pm 0.1}$
Asterix $\mathbf{1.3}\scriptstyle{\pm 0.1}$ $7.2\scriptstyle{\pm 0.2}$
BankHeist $\mathbf{2.9}\scriptstyle{\pm 0.3}$ $8.9\scriptstyle{\pm 0.6}$
BattleZone $\mathbf{2.2}\scriptstyle{\pm 0.3}$ $9.6\scriptstyle{\pm 0.6}$
ChopperCommand $\mathbf{3.4}\scriptstyle{\pm 0.4}$ $9.0\scriptstyle{\pm 0.7}$
CrazyClimber $\mathbf{2.7}\scriptstyle{\pm 0.1}$ $9.1\scriptstyle{\pm 0.4}$
DemonAttack $\mathbf{1.1}\scriptstyle{\pm 0.1}$ $6.8\scriptstyle{\pm 0.8}$
Freeway $\mathbf{1.0}\scriptstyle{\pm 0.0}$ $6.1\scriptstyle{\pm 0.2}$
Frostbite $\mathbf{2.2}\scriptstyle{\pm 0.4}$ $10.9\scriptstyle{\pm 0.8}$
Gopher $\mathbf{3.2}\scriptstyle{\pm 0.6}$ $8.1\scriptstyle{\pm 0.8}$
Hero $\mathbf{2.5}\scriptstyle{\pm 0.4}$ $9.9\scriptstyle{\pm 0.6}$
Jamesbond $\mathbf{2.2}\scriptstyle{\pm 0.3}$ $9.1\scriptstyle{\pm 0.5}$
Kangaroo $\mathbf{2.0}\scriptstyle{\pm 0.2}$ $8.6\scriptstyle{\pm 0.8}$
Krull $\mathbf{1.8}\scriptstyle{\pm 0.1}$ $7.7\scriptstyle{\pm 0.3}$
KungFuMaster $\mathbf{1.3}\scriptstyle{\pm 0.1}$ $7.6\scriptstyle{\pm 0.7}$
PrivateEye $\mathbf{1.0}\scriptstyle{\pm 0.1}$ $5.8\scriptstyle{\pm 0.5}$
RoadRunner $\mathbf{1.5}\scriptstyle{\pm 0.2}$ $9.0\scriptstyle{\pm 0.3}$
UpNDown $\mathbf{1.4}\scriptstyle{\pm 0.1}$ $7.2\scriptstyle{\pm 0.4}$
Pong $\mathbf{1.0}\scriptstyle{\pm 0.1}$ $4.0\scriptstyle{\pm 0.5}$
MsPacman $\mathbf{1.4}\scriptstyle{\pm 0.2}$ $6.9\scriptstyle{\pm 0.3}$
Qbert $\mathbf{1.3}\scriptstyle{\pm 0.1}$ $7.0\scriptstyle{\pm 0.3}$
Seaquest $\mathbf{1.9}\scriptstyle{\pm 0.4}$ $10.1\scriptstyle{\pm 0.5}$
Boxing $\mathbf{1.1}\scriptstyle{\pm 0.0}$ $6.6\scriptstyle{\pm 0.1}$
Breakout $\mathbf{3.0}\scriptstyle{\pm 0.8}$ $4.9\scriptstyle{\pm 1.8}$
Table 6: Average wall-time(hours) of ReZero-M vs. MuZero on 26 Atari game environments. The time represents the average total wall-clock time to 100k environment steps. Mean and standard deviation over 5 runs.
Figure 7 displays the performance over environment interaction steps of the ReZero-M algorithm compared with the original MuZero algorithm across six representative Atari environments and two board games. We can find that ReZero-M obtained similar sample efficiency than MuZero on the most tasks.